OpenStudy (anonymous):

A function \[f(x)=log(x), x > 0\] is given. Come with a good guess what \[d^n/dx^n\] is, when n is a natural number. Proof that your guess is correct.

OpenStudy (badhi):

It would be easier if you can mention where you are stuck at ;)

OpenStudy (anonymous):

f' = 1/x f'' = -1/x^2 f''' = 1*2/x^3 Hence, f^n = (-1)^n-1 * n-1!/x^n